OpenStudy (erniellyd28):

a certain type of exercise can help a person burn fat and lose weight.each additional minute of this exercise burns an extra 20 calories per day.write a linear function that models the number of calories burned each day by x minutes of this exercise.by burning an additional 3500 calories, a person can lose 1 pond of fat.how many minutes of this particular exercise are needed to burn 1 pound of fat in 15 days? ( write a linear function that models the number of calories burned each day by x minutes of this exercise)

OpenStudy (thereisnousername):

x minutes______________calories PER DAY 0______________________0 1_____________________30 2_____________________2*30 3_____________________3*30 . . n_____________________x*30 f(x)=30x F(x) is how many calories per day for x minutes of the exercise. Now is F(x) in CALORIES per DAY which depends on x. Given also 3500 calories = 1 pound of fat, (1/3500)*f(x) gives how many POUNDS per DAY. How can this be formed into a uniform rates relationship? R*T=M burn rate, time in days, mass of fat in pounds. (1/3500)*f(x)*t=m changing the names of the variables for distinction; Let t=10, m=1; (1/3500)f(x)*10=1 REMEMBER that f(x)=30x and x was minutes of exercise in the any one day. (1/3500)*30x*10=1 now just solve for x

OpenStudy (erniellyd28):

so the answer is 1/3500*30x*10=1?